# Matter-of-Fact: A Benchmark for Verifying the Feasibility of Literature-Supported Claims in Materials Science

### v1

@@ 4.3 Results
Feasibility Assessment Results:
The performance of all models when evaluated in the feasibility assessment mode is shown in Table 2. Performance across all models ranges from 0.58 to 0.72, with the models that use the smallest (and least expensive) base model (GPT-4o-mini) generally performing about 5 percent lower than the two more performant (and more costly) base models, o4-mini and Claude Sonnet 3.7. Across models, adding features (such as in-context learning, reflection, RAG over SemanticScholar, or Code Generation) generally provides modest performance improvements, or does not improve performance over the Chain-of-Thought baseline, highlighting the difficulty of this task when using conventional solution methods, and its suitability as a challenge task. When examining performance broken down by category, we observe that while the overall performance of a given base model is similar with different features, some models are more performant at identifying true/feasible claims than they are at identifying false/infeasible claims, and vice versa. All models perform better at assessing the feasibility of qualitative claims than quantitative claims, with this difference between 11% and 19% across all models, potentially a result of quantitative claims requiring the ability (through code or other means) of verifying the feasibility of specific numerical values present in the claims. In line with this reasoning, claims that are based on experiments or code/simulations consistently achieve the lowest performance, while those based on theoretical results are next-highest, with integrative claims achieving the highest performance.
@@
Table 4: Knowledge contamination analysis of base models. In this analysis, performance of the Chain-of-Thought + 20-Shot ICL + Reflection model is shown for claims from papers that were authored before or after a given model’s advertised knowledge cutoff date. Given the temporal nature of the dataset, all 8.4k claims across train, development, and test sets were included. All models show almost identical performance (±$\pm 1\%$) when tested on claims from papers before or after their knowledge cutoff date, suggesting that knowledge contamination does not play a significant role in performance.
Claim Verification Results:
The performance of models when evaluated in the claim verification mode is shown in Table 3. In this mode the models have no temporal restrictions, and may use knowledge from the source paper, or papers authored after the source paper (including those that may cite the source paper) as evidence to perform the claim verification task. These experiments serve two purposes. First, they identify an effective ceiling of how well a given base model can perform even when provided with the original source article used to create a claim, with o4-mini and Claude Sonnet 3.7 capable of achieving nearly a 100% ceiling performance, while gpt-4o-mini has more modest performance ceiling between 0.71 and 0.76. Second, these models serve as a consistency evaluation for the claim generation protocol, emphasizing that when strong models are asked to verify the labels of these automatically generated claims, they nearly always agree with the gold label. Further emphasizing this is the domain expert performance, who (when provided with the original source article), agreed with the LLM-generated label for 99% of claims after resolving disagreements.
@@ 5 Discussion
Base-Model Contamination:
@@
A central part of the framing of our feasibility assessment task as a temporally-filtered claim verification task is that it requires models to have a minimum of contamination with knowledge beyond a given claim’s knowledge cutoff date. While it is possible that techniques such as model editing and machine unlearning Bourtoule et al. (2021); Tarun et al. (2023); Liu et al. (2025) may eventually allow the knowledge in a base model to be temporally filtered to minimize this contamination, this may have limited success in current forms Lynch et al. (2024); Deeb and Roger (2024); Du et al. (2024). Instead, here we aim to measure how much of the current model performance is likely due to model contamination (from, for example, the base model being trained on the source articles used to generate the claims). To measure this, we examine each base model’s performance for claims extracted from papers before and after the base model’s advertised training data knowledge cut-off dates. The results, shown in Table 4, show that the performance of the base models on claims from papers authored after their knowledge cutoff is nearly identical to the performance on claims authored by papers that are before the knowledge cutoff date. This empirically suggests that the performance of current base models on the feasibility assessment task is not due to model contamination, but due to other properties, such as their capacity for reasoning.
Pragmatic Ceiling Performance:
@@
While we empirically show that the feasibility of many claims can be assessed using inexpensive means, the models we demonstrate are far from achieving perfect performance on this task. Pragmatically, a model that achieves near 100% performance would be able to (with near perfect accuracy) determine whether claims are likely to be feasible or infeasible through some combination of literature search, inexpensive code-based experimentation, world modeling, and other means. Acheiving 100% performance is likely impractical, as many scientific claims can only be verified with empirical work, and not with literature search or simulation, particularly for those (most impactful) scientific results that are surprising because they run counter to expectations. That being said, even though effective ceiling performance on feasibility assessment tasks is likely to be less than 100%, increasing model performance on this task even a modest amount can have practical utility for improving the efficiency of discovery systems. As we show in Appendix A, for a hypothetical hypothesis generation system where 1% of the hypotheses are true, the performance of our current-best model could potentially allow discovering 60% of the true hypotheses while reducing experiment costs by 80% – a large overall budget reduction, at the cost of reducing the recall of finding true hypotheses by approximately 40%.
@@ 6 Conclusion
Temporal Filtering for Prediction:
Temporal datasets offer the opportunity to construct prediction datasets for high-impact domains (e.g. Luo et al., 2018, link prediction for cancer biology) where the knowledge a system is predicting is potentially beyond current human knowledge, and for which gold labels are infeasible to construct. Temporal filtering assumes well-controlled models that have not been contaminated with data past their temporal filtering date. In this work we characterize the contamination rate of our base language models, and this analysis suggests that data contamination either does not exist, or is not a significant factor driving current performance. That being said, users of this benchmark should characterize the performance of novel base models to characterize how much data contamination may play a role.
Limits of Code-based Experimentation:
Pragmatically, to be useful for filtering scientific hypotheses, feasibility assessment methods must be able to perform well at scale. This necessitates that any pilot experiments (including code-based simulations) must be fast and inexpensive to run, otherwise the feasibility assessment step may be impractically expensive to provide overall cost savings. That being said, different applications and end-users may have varying preferred cost/performance points, and we encourage reporting performance as a function of overall cost (as we have done in this work) to help accurately assess the cost vs benefit of proposed models. It is our hope that providing a large-scale benchmark that necessitates developing inexpensive feasibility assessment methods will help facilitate innovation in this direction.
@@ Acknowledgments
This research was developed with funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300) to PJ at the University of Arizona. The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government. PJ has an outside interest in the Allen Institute for Artificial Intelligence. This interest has been disclosed to the University of Arizona and reviewed in accordance with its conflict of interest policies. We thank the members of the DARPA Scientific Feasibilty (SciFy) program for thoughtful discussions.
@@ Appendix A Utility for Hypothesis Filtering
Feasibility assessment has utility in impactful tasks such as (semi-automated) scientific discovery, particularly in the context of hypothesis generation. Hypothesis generation systems (e.g Lu et al., 2024; Jansen et al., 2025; O’Neill et al., 2025) have the capacity to generate an impractically large number of possible hypotheses (framed as claims) that one could test, and as a result running all their proposed experiments is costly (at best) and intractible (at worst). Coupling hypothesis generation with feasibility assessment would allow filtering out hypotheses that are unlikely to be feasible – i.e. yield experimental results that support the hypothesis – and ultimately increase the efficiency of scientific discovery systems in terms of the number of positive discoveries that can be made on a given budget. In automated hypothesis generation where overall likelihood of a hypothesis yielding positive results is low, increasing efficiency is dominated by correctly identifying (and filtering) infeasible hypotheses/claims. Figure 3 shows a plot of experiment efficiency (in terms of the reduction in the number of experiments that would need to be run) for hypothetical hypothesis generation systems that have different rates of generating true hypotheses, with the performance of the best-performing model (RAG (Semantic Scholar) using o4-mini) highlighted. For a hypothetical hypothesis generation system where 1% of its hypotheses are true, this model would reduce the number of experiments needed to be run by approximately 80%, while still discovering 60% of the true hypotheses. This highlights that even systems with middle performance can have practical utility (in terms of cost savings) when coupled with scientific discovery systems.

### v2

@@ 4.3 Results
Feasibility Assessment Results: The performance of all models when evaluated in the feasibility assessment mode is shown in Table 2. Performance across all models ranges from 0.58 to 0.72, with the models that use the smallest (and least expensive) base model (GPT-4o-mini) generally performing about 5 percent lower than the two more performant (and more costly) base models, o4-mini and Claude Sonnet 3.7. Across models, adding features (such as in-context learning, reflection, RAG over SemanticScholar, or Code Generation) generally provides modest performance improvements, or does not improve performance over the Chain-of-Thought baseline, highlighting the difficulty of this task when using conventional solution methods, and its suitability as a challenge task. When examining performance broken down by category, we observe that while the overall performance of a given base model is similar with different features, some models are more performant at identifying true/feasible claims than they are at identifying false/infeasible claims, and vice versa. All models perform better at assessing the feasibility of qualitative claims than quantitative claims, with this difference between 11% and 19% across all models, potentially a result of quantitative claims requiring the ability (through code or other means) of verifying the feasibility of specific numerical values present in the claims. In line with this reasoning, claims that are based on experiments or code/simulations consistently achieve the lowest performance, while those based on theoretical results are next-highest, with integrative claims achieving the highest performance.
@@
Table 4: Knowledge contamination analysis of base models. In this analysis, performance of the Chain-of-Thought + 20-Shot ICL + Reflection model is shown for claims from papers that were authored before or after a given model’s advertised knowledge cutoff date. Given the temporal nature of the dataset, all 8.4k claims across train, development, and test sets were included. All models show almost identical performance ($\pm 1\%$) when tested on claims from papers before or after their knowledge cutoff date, suggesting that knowledge contamination does not play a significant role in performance.
Claim Verification Results: The performance of models when evaluated in the claim verification mode is shown in Table 3. In this mode the models have no temporal restrictions, and may use knowledge from the source paper, or papers authored after the source paper (including those that may cite the source paper) as evidence to perform the claim verification task. These experiments serve two purposes. First, they identify an effective ceiling of how well a given base model can perform even when provided with the original source article used to create a claim, with o4-mini and Claude Sonnet 3.7 capable of achieving nearly a 100% ceiling performance, while gpt-4o-mini has more modest performance ceiling between 0.71 and 0.76. Second, these models serve as a consistency evaluation for the claim generation protocol, emphasizing that when strong models are asked to verify the labels of these automatically generated claims, they nearly always agree with the gold label. Further emphasizing this is the domain expert performance, who (when provided with the original source article), agreed with the LLM-generated label for 99% of claims after resolving disagreements.
@@ 5 Discussion
5.1 Controlling Potential Confounds
@@
Base-Model Contamination: A central part of the framing of our feasibility assessment task as a temporally-filtered claim verification task is that it requires models to have a minimum of contamination with knowledge beyond a given claim’s knowledge cutoff date. While it is possible that techniques such as model editing and machine unlearning Bourtoule et al. (2021); Tarun et al. (2023); Liu et al. (2025) may eventually allow the knowledge in a base model to be temporally filtered to minimize this contamination, this may have limited success in current forms Lynch et al. (2024); Deeb and Roger (2024); Du et al. (2024). Instead, here we aim to measure how much of the current model performance is likely due to model contamination (from, for example, the base model being trained on the source articles used to generate the claims). To measure this, we examine each base model’s performance for claims extracted from papers before and after the base model’s advertised training data knowledge cut-off dates. The results, shown in Table 4, show that the performance of the base models on claims from papers authored after their knowledge cutoff is nearly identical to the performance on claims authored by papers that are before the knowledge cutoff date. This empirically suggests that the performance of current base models on the feasibility assessment task is not due to model contamination, but due to other properties, such as their capacity for reasoning.
Collecting Claims from Scientific Papers: In this work, the claims we use for the feasibility assessment task (and the associated hypotheses underlying those claims) are collected from scientific articles. It is common that the narrative presented in an article differs from the actual scientific process that was undertaken, which typically includes failed experiments, promising initial directions that turned out to be dead ends, and other complexities that are often left out when writing a paper. As such, the claims (and associated hypotheses) extracted from papers may have a different (and more polished) character than those one naturally explores during the scientific discovery process. While it is currently difficult to quantify this potential difference, we wish to acknowledge that it may exist, and that this may ultimately affect transfer performance in models trained or evaluated on literature-derived claims (like those in Matter-of-Fact) to real-world discovery scenarios.
5.2 Performance Characterization
@@
Pragmatic Ceiling Performance: While we empirically show that the feasibility of many claims can be assessed using inexpensive means, the models we demonstrate are far from achieving perfect performance on this task. Pragmatically, a model that achieves near 100% performance would be able to (with near perfect accuracy) determine whether claims are likely to be feasible or infeasible through some combination of literature search, inexpensive code-based experimentation, world modeling, and other means. Achieving 100% performance is likely impractical, as many scientific claims can only be verified with empirical work, and not with literature search or simulation, particularly for those (most impactful) scientific results that are surprising because they run counter to expectations. That being said, even though effective ceiling performance on feasibility assessment tasks is likely to be less than 100%, increasing model performance on this task even a modest amount can have practical utility for improving the efficiency of discovery systems. As we show in Appendix A, for a hypothetical hypothesis generation system where 1% of the hypotheses are true, the performance of our current-best model could potentially allow discovering 60% of the true hypotheses while reducing experiment costs by 80% – a large overall budget reduction, at the cost of reducing the recall of finding true hypotheses by approximately 40%.
Limitations in Baseline Models (Retrieval): It is entirely plausible that assessing the feasibility of complex state-of-the-art scientific claims would require integrating knowledge that comes from more than one scientific article. While our retrieval-augmented-generation system presents the top-K paper snippets to the model making the feasibility assessment, in our baseline system these snippets come from a single search query, and multiple search queries may be required to collect different types of evidence that, together, could be integrated to improve the feasibility assessment. An initial pilot system that we constructed that iteratively allowed up to 5 rounds of evidence collection (each with a different query) before making an assessment did not appear to improve feasibility performance, suggesting integrating this knowledge into improved feasibility task performance is non-trivial. Similarly, it is plausible that building a specialized domain-specific corpus of supporting scientific articles may improve the utility of the retrieved knowledge, and increase task performance.
Limitations in Baseline Models (Code): Due to the current high estimated cost in running code generation systems, and the large size of the Matter-of-Fact test set, our code generation baseline was run in a highly limited form (i.e. short runtime, no debug iterations) that is best considered an approximate measure of zero-shot code generation performance on this feasibility assessment task, without the benefits of debugging iterations or the long experiment runtimes that code-based solutions to this task would almost certainly require. Some of these limitations are pragmatic, like cost and runtime, and are rapidly reduced as (for example) open language models for code generation that can be run on local hardware begin to approach paid API-based model performance. However, we ultimately believe that progress in code-based experimentation will require building systems that integrate materials science domain tools, software, and databases. We did not include any materials-science specific tooling in CodeScientist’s code retrieval library and required it to instead rely on the base model weights to implement these complex tool interfaces. It is highly likely that near-term improvement on this task will require a degree of manually integrating this tooling, just as other scientific agents (such as Biome Huang et al. (2025) in the biomedical domain) are currently using hand-build interfaces to domain-specific tools and databases to support their discovery tasks.
@@ 6 Conclusion
Temporal Filtering for Prediction: Temporal datasets that use the notion of backtesting offer the opportunity to construct prediction datasets for high-impact domains (e.g. Luo et al., 2018, link prediction for cancer biology) where the knowledge a system is predicting is potentially beyond current human knowledge, and for which gold labels are infeasible to construct. Temporal filtering assumes well-controlled models that have not been contaminated with data past their temporal filtering date. In this work we characterize the contamination rate of our base language models, and this analysis suggests that data contamination either does not exist, or is not a significant factor driving current performance. That being said, users of this benchmark should characterize the performance of novel base models to characterize how much data contamination may play a role.
Cost-Benefit Analyses: Pragmatically, to be useful for filtering scientific hypotheses, feasibility assessment methods must be able to perform well at scale. This necessitates that any pilot experiments (including code-based simulations) must be fast and inexpensive to run, otherwise the feasibility assessment step may be impractically expensive to provide overall cost savings. That being said, different applications and end-users may have varying preferred cost/performance points, and we encourage reporting performance as a function of overall cost (as we have done in this work) to help accurately assess the cost vs benefit of proposed models. It is our hope that providing a large-scale benchmark that necessitates developing inexpensive feasibility assessment methods will help facilitate innovation in this direction.
@@ Acknowledgments
This research was developed with funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300) to PJ at the University of Arizona. The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government. PJ has an outside interest in the Allen Institute for Artificial Intelligence. This interest has been disclosed to the University of Arizona and reviewed in accordance with its conflict of interest policies. We thank the members of the DARPA Scientific Feasibility (SciFy) program for thoughtful discussions. We also wish to thank the anonymous reviewers for their helpful comments, and their particular observation that literature-derived claims may be more polished than those present in earlier stages of the discovery process.
@@ Appendix A Utility for Hypothesis Filtering
Feasibility assessment has utility in impactful tasks such as (semi-automated) scientific discovery, particularly in the context of hypothesis generation. Hypothesis generation systems (e.g Lu et al., 2024; Jansen et al., 2025; O’Neill et al., 2025) have the capacity to generate an impractically large number of possible hypotheses (framed as claims) that one could test, and as a result running all their proposed experiments is costly (at best) and intractable (at worst). Coupling hypothesis generation with feasibility assessment would allow filtering out hypotheses that are unlikely to be feasible – i.e. yield experimental results that support the hypothesis – and ultimately increase the efficiency of scientific discovery systems in terms of the number of positive discoveries that can be made on a given budget. In automated hypothesis generation where overall likelihood of a hypothesis yielding positive results is low, increasing efficiency is dominated by correctly identifying (and filtering) infeasible hypotheses/claims. Figure 3 shows a plot of experiment efficiency (in terms of the reduction in the number of experiments that would need to be run) for hypothetical hypothesis generation systems that have different rates of generating true hypotheses, with the performance of the best-performing model (RAG (Semantic Scholar) using o4-mini) highlighted. For a hypothetical hypothesis generation system where 1% of its hypotheses are true, this model would reduce the number of experiments needed to be run by approximately 80%, while still discovering 60% of the true hypotheses. This highlights that even systems with middle performance can have practical utility (in terms of cost savings) when coupled with scientific discovery systems.